OpenStudy (anonymous):

Solve the system of linear equations and check any solution algebraically. 2x +2z = 6 5x +3y = 11 3y -4z = 1

OpenStudy (anonymous):

You know about operations inside a matrix?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

also using elimination method.., do you know???

OpenStudy (anonymous):

Try substracting line 3 to line 2, what does line 2 become?

OpenStudy (anonymous):

hey get z in terms of x from 1st equation, y in terms of x from 2nd equation. subtite both in third equation ,you will get a value for x. from then calculate z and y.

OpenStudy (anonymous):

@theloser it can be done that way, but it'll probably be worth a 0 in an exam on Linear Algebra. They want to pave the way for more complicated problems using matrices.

OpenStudy (anonymous):

Line 2 would become 5x+0+4z=10 Do you see what I've done?

OpenStudy (anonymous):

@AlexUL i got -5x +3y = -10 but i clearly did something wrong

OpenStudy (anonymous):

@AlexUL never mind i got that too now

OpenStudy (anonymous):

2x +2z = 6 |x5| 5x +3y = 11 |x2| --------------- - 10x +10z = 30 10x +6y = 22 -------------- - 10z - 6y = -2 -6y + 10z = -2 |x1| 3y -4z = 1 |x2| -6y + 10z = -2 6y -8z = 2 -------------- + 0 + 2z = 0 z = 0 so 10z -6 = -2 0 -6y = -2 -6y = -2 y = 1/3 10x +6y = 22 10x + 6 (1/3) = 22 10x + 2 = 22 10x = 22-2 10x = 20 x = 2

OpenStudy (anonymous):

ok, good! :P Now, how can you simplify even more line 2, so you end up with only x or z?

OpenStudy (anonymous):

x = 2 y = 1/3 z = 0

OpenStudy (anonymous):

@gerryliyana the answer in the back of the book says (-2, 7, 5)

OpenStudy (anonymous):

sorry i'm wrong

OpenStudy (anonymous):

@gerryliyana it's okay :)

OpenStudy (anonymous):

@AlexUL who said that have to be solved by matrices only. btw here the coeff' matrix will have 1 zero in all columns and rows. so the method i have told works better. if the coefficient matrix dint have any zeros, then taking inverse and all that stuff works better.

OpenStudy (anonymous):

x=-2,y=7,z=5

OpenStudy (anonymous):

@pkjha3105 how'd u do that?

OpenStudy (anonymous):

i hv solvd in rough...doesnt it satisy the eq?

OpenStudy (anonymous):

@pkjha3105 that's the correct answer but i wanna know how to solve it

OpenStudy (anonymous):

rewrite 1st equation as z = 3-x 2nd as \[y = \frac{ 11-5x }{ 3 }\] put and y and z values in third equation

OpenStudy (anonymous):

solve by matrix methrd...put all constant in one 3 by 3 matrix,multiply it by matrix of varible and equate to matrix,wch contains the value of rhs....then find the inverse of matrix containg constant n multiply by rhs....u will get the value of x,y,z...

OpenStudy (anonymous):

2x +2z = 6 |x5| 5x +3y = 11 |x2| --------------- - 10x +10z = 30 10x +6y = 22 -------------- - 10z - 6y = 8 -6y + 10z = 8 |x1| 3y -4z = 1 |x2| -6y + 10z = 8 6y -8z = 2 -------------- + 0 + 2z = 10 z = 10/2 z = 5 so 10z -6 = 8 10 (5) -6y = 8 -6y = 8 -50 y = -42/-6 y =7 5x +3y = 11 5x + 3(7) = 11 5x + 21 = 11 5x = 11-21 x = -10/5 x =-2

OpenStudy (anonymous):

how about that @lawls ??

OpenStudy (anonymous):

the answer is x = -2 y = 7 z =5 take it easy.,

OpenStudy (anonymous):

@theloser so 3((11-5x)/3) - 4(3-x) = 1

OpenStudy (anonymous):

heiii @lawls look at my LAST solution!

OpenStudy (anonymous):

@gerryliyana good job u got the answer. i'm gonna see how u did it now.

OpenStudy (anonymous):

@lawls thank you :)

OpenStudy (anonymous):

@theloser The question says "check any answer algebraically", I think this means you have to find the answer in another way, and if lawls is studying systems of linear equations and has written the system as a matrix, I assume you need to work it out with matrices operations

OpenStudy (anonymous):

@lawls .. exactly, from that you will get x= -2, subtitute in z and y exprestion which are in terms of x...

OpenStudy (anonymous):

@AlexUL , no check answer algebraically , i think would mean, keep the point in all there equations. and if it is correct i will satisfy them

OpenStudy (anonymous):

it is preferable to solve these eq. by matrix methord only....find the inverse of matrix containing const. and multiply by the matrix of rhs....x,y,z...can easily b found by this methord...

OpenStudy (anonymous):

@pkjha3105 & @AlexUL my teacher hasn't taught us about matrices yet. in class he solved these types of questions in @gerryliyana way so i'm gonna use that, but thanks so much though.

OpenStudy (anonymous):

@theloser i got the answer in your way as well. it was very simple. thanks :)

OpenStudy (anonymous):

ur welcome @lawls good job! :)